OpenStudy (samhang):

given vector a=3i+4j, vector b=8i+6j and vector c=2i+kj find the value of k, here vector c makes equal angle with vectors a and b.

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle {\bf a}=\langle 3,4\rangle}\) \(\color{black}{\displaystyle {\bf b}=\langle 3m,4m\rangle~~~{\rm such~that}\,~k>0}\) \(\color{black}{\displaystyle {\bf c}=\langle 2,k\rangle}\) \(\color{blue}{\displaystyle {\bf a}\cdot {\bf c}=\left|{\bf a}\right|\left|{\bf c}\right|\cos \theta_1 }\) \(\color{blue}{\displaystyle 6+4k=5\sqrt{k^2+4}\cos\theta_1 }\) \(\color{green}{\displaystyle {\bf b}\cdot {\bf c}=\left|{\bf b}\right|\left|{\bf c}\right|\cos \theta_2 }\) \(\color{green}{\displaystyle 6m+4mk=5m\sqrt{k^2+4}\cos\theta_2}\) \(\color{blue}{\displaystyle \cos\theta_1= \frac{6+4k}{5\sqrt{k^2+4}} }\) \(\color{green}{\displaystyle \cos\theta_2= \frac{6m+4mk}{5m\sqrt{k^2+4}}=\frac{6+4k}{5\sqrt{k^2+4}} }\)

OpenStudy (solomonzelman):

So, you can see that in the above case (which is the same as yours, but a bit more abstract), regardless of the \(m\), the angles are same.

OpenStudy (solomonzelman):

This is because \(\bf a\) and \(\bf b\) are in the same direction, so, regardless of their magnitudes (as long as the corresponding components of \(\bf a\) and \(\bf b\) are proportional), they will always make the same angle with any other vector \(\bf c\), no matter what this vector \(\bf c\) is (well, asides from \(\bf 0\)).

OpenStudy (solomonzelman):

Oh, I am sorry, they are not proportional .... it is 8,6 not 6,8 :(

OpenStudy (solomonzelman):

Well, you can follow the same process with your vectors: \(\color{black}{\displaystyle {\bf a}=\langle 3,4\rangle}\) \(\color{black}{\displaystyle {\bf b}=\langle 8,6\rangle~~~{\rm such~that}\,~k>0}\) \(\color{black}{\displaystyle {\bf c}=\langle 2,k\rangle}\) \(\color{blue}{\displaystyle {\bf a}\cdot {\bf c}=\left|{\bf a}\right|\left|{\bf c}\right|\cos \theta_1 }\) \(\color{blue}{\displaystyle 6+4k=5\sqrt{k^2+4}\cos\theta_1 }\) \(\color{green}{\displaystyle {\bf b}\cdot {\bf c}=\left|{\bf b}\right|\left|{\bf c}\right|\cos \theta_2 }\) \(\color{green}{\displaystyle 16+6k=10\sqrt{k^2+4}\cos\theta_2}\) \(\color{blue}{\displaystyle \cos\theta_1= \frac{6+4k}{5\sqrt{k^2+4}} }\) \(\color{green}{\displaystyle \cos\theta_2= \frac{16+6k}{10\sqrt{k^2+4}} }\)

OpenStudy (solomonzelman):

Then, set \(\color{black}{\displaystyle \cos\theta_1= \cos\theta_2 }\). \(\color{black}{\displaystyle \frac{6+4k}{5\sqrt{k^2+4}} =\frac{16+6k}{10\sqrt{k^2+4}}}\)

OpenStudy (samhang):

let me try

OpenStudy (samhang):

I got k = 2...

OpenStudy (samhang):

Thanks a lot

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle \frac{6+4k}{5\sqrt{k^2+4}} =\frac{16+6k}{10\sqrt{k^2+4}}}\) \(\color{black}{\displaystyle 6+4k=8+3k}\) \(\color{black}{\displaystyle k=2}\)

OpenStudy (solomonzelman):

Correct ... You welcome!